OpenStudy (anonymous):

Help Please<3 Activity 1. 2. a. What is the common ratio of the sequence 80, 20, 5, 5/4, ...? b. What are the next three terms in the sequence? 3. a. What is the common ratio of the sequence 2, -6, 18, -54, ...? b. What are the next three terms in the sequence? Thank You <3

OpenStudy (anonymous):

(a)common ratio is r = t2 / t1 so r=20/80=1/4 u can calculate the next 3 terms using common ratio

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

what's so confusing...... see common ratio is same for all terms here.......